# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ Appendix A Discrete Differential Calculus for Abelian Fields
$0\to\mathcal{C}^{0}\xrightarrow{\mathrm{d}_{0}}\mathcal{C}^{1}\xrightarrow{\mathrm{d}_{1}}\cdots\xrightarrow{\mathrm{d}_{D-1}}\mathcal{C}^{D}\to 0,$ (133)
as
$H^{k}:=\mathrm{ker}\,\mathrm{d}_{k}/\mathrm{im}\,\mathrm{d}_{k-1}.$ (134)
Here $\mathrm{ker}\,\mathrm{d}_{k}$ is the set of locally-constant $k$-forms (called closed forms), and $\mathrm{im}\,\mathrm{d}_{k-1}$ is the set of $k$-forms which are gradients of $(k-1)$-forms (called exact forms). Every exact form is closed, but there can exist closed forms which are not exact, which are classified by cohomology. Cohomology classes are equivalence classes of closed forms which differ by an exact form, just as homology classes are equivalence classes of closed surfaces which differ by a boundary. Note that exact forms integrate to zero by Stoke’s theorem on any closed surface, while closed forms need only integrate to zero on closed surfaces that are boundaries. Forms with non-trivial cohomology class integrate non-trivially on non-contractible surfaces, defining a pairing between cohomology and homology classes.
@@ Appendix B U(1) Bulk Phase Diagram: Limits, Symmetries and Monte Carlo
Here we provide a brief review of the bulk phase diagram of the 4D $U(1)$ Abelian-Higgs model, described by the action Eq. 8, the Hamiltonian Eq. 10, and dual action Eq. 108, by considering the various limits and undertaking numerical Monte Carlo simulation in the action formulation. The phase diagram is sketched in Fig. 1, which conveniently summarizes the discussion that follows. In particular, it is well-known that there are only two distinct phases [12], the Coulomb phase and the Higgs-confined phase. According to Eq. 102, this is a theory of electric strings terminating on electric point charges. The dual magnetic description is in terms of magnetic strings terminating on magnetic point charges (vortices and monopoles of the Higgs and gauge fields, respectively). The basic structure of the phase diagram can be deduced by consider each of the four limits.
Pure gauge limit ($\kappa\ll 1$): In the limit $\kappa=0$ the gap of the electric point charges diverges, and the theory reduces to 4D $U(1)$ gauge theory. In this limit, the system has a global 1-form symmetry, $\cramped{A\to A+\lambda}$ with $\cramped{\mathrm{d}\lambda=0}$, called electric symmetry. This symmetry yields corresponds to electric strings forming closed loops. This theory has two phases, a confined phase at small $\beta$ (strong coupling), where the system is gapped and electric strings cost energy proportional to their length; and a deconfined phase at large $\beta$ (weak coupling), where the electric strings condense, the electric symmetry is spontaneously broken, and the system has a gapless photon excitation. Turning on a small $\kappa$ explicitly breaks the 1-form symmetry by introducing gapped electric point charges at which open electric strings end. Because the charges are strongly gapped, qualitatively speaking we expect the 1-form symmetry to re-emerge at low energies below the charge gap, thus allowing the deconfined phase to extend to finite $\kappa$.
@@
Strong coupling limit ($\beta=0$): In this limit the curvature of the gauge field $A$ is not penalized, and we can say that the magnetic monopoles are maximally proliferated. If we fix to unitary gauge ($\theta=\text{const}.$), the action becomes $-\kappa\sum_{\ell}\cos(A_{\ell})$, reducing to a set of completely disconnected link variables. Thus the system is in a trivial phase for all $\kappa$, and there is no bulk phase transition on this line.
Infinite Higgs coupling limit ($\kappa\to\infty$): In this limit we have the constraint $A=\mathrm{d}\theta$. This implies that the electric charge creation and annihilation operators, Eq. 14, act as the identity, and the system can be described as an electric condensate. The bulk has no dynamics and is completely frozen, which can be seen by fixing to unitary gauge ($\theta=\text{const}.$), in which $A=0$ and the Higgs field is frozen. There is therefore no bulk phase transition along this line. For large but finite $\kappa$ the bulk action can be roughly understood as the proca-type action, Eq. 23, describing a massive 1-form field.
This completes the general outline of the phase diagram in the vicinity of the edges in Fig. 1. The only question that remains is how the two transitions at small $\kappa$ and large $\beta$ reach each other to separate the confined phase (in which the electric and magnetic symmetries are emergent and spontaneously broken) from the Higgs-confined phase. Figure 12 shows results from Monte Carlo simulations measuring the average Wilson plaquette and Wilson link, which fills in the remainder of the phase diagram schematically shown in Fig. 1. The two transitions extend as first-order transition lines and meet at a triple point in the vicinity of $\beta\sim 1.0$ an $\kappa\sim 0.4$. A third first order line extends from the this triple point towards smaller $\beta$ and larger $\kappa$ which ends at a critical endpoint. We show evidence for this first-order line in Fig. 12(c).

### v2

@@ Appendix A Discrete Differential Calculus for Abelian Fields
$0\to\mathcal{C}^{0}\xrightarrow{\mathrm{d}_{0}}\mathcal{C}^{1}\xrightarrow{\mathrm{d}_{1}}\cdots\xrightarrow{\mathrm{d}_{D-1}}\mathcal{C}^{D}\to 0,$ (A.15)
as $H^{k}:=\mathrm{ker}\,\mathrm{d}_{k}/\mathrm{im}\,\mathrm{d}_{k-1}$. Here $\mathrm{ker}\,\mathrm{d}_{k}$ is the set of locally-constant $k$-forms (called closed forms), and $\mathrm{im}\,\mathrm{d}_{k-1}$ is the set of $k$-forms which are gradients of (k1)(k-1)-forms (called exact forms). Every exact form is closed, but there can exist closed forms which are not exact, which are classified by cohomology. Cohomology classes are equivalence classes of closed forms which differ by an exact form, just as homology classes are equivalence classes of closed surfaces which differ by a boundary. Note that exact forms integrate to zero by Stoke’s theorem on any closed surface, while closed forms need only integrate to zero on closed surfaces that are boundaries. Forms with non-trivial cohomology class integrate non-trivially on non-contractible surfaces, defining a pairing between cohomology and homology classes.
@@ Appendix B U(1) Bulk Phase Diagram: Limits, Symmetries and Monte Carlo
Here we provide a brief review of the bulk phase diagram of the 4D $U(1)$ Abelian-Higgs model, described by the action Eq.˜8, the Hamiltonian Eq.˜10, and dual action Eq.˜110, by considering the various limits and undertaking numerical Monte Carlo simulation in the action formulation. The phase diagram is sketched in Fig.˜1, which conveniently summarizes the discussion that follows. In particular, it is well-known that there are only two distinct phases [14], the Coulomb phase and the Higgs-confined phase. According to Eq.˜104, this is a theory of electric strings terminating on electric point charges. The dual magnetic description is in terms of magnetic strings terminating on magnetic point charges (vortices and monopoles of the Higgs and gauge fields, respectively). The basic structure of the phase diagram can be deduced by consider each of the following four limits.
Pure gauge limit ($\kappa\ll 1$): In the limit $\kappa=0$ the gap of the electric point charges diverges, and the theory reduces to 4D $U(1)$ gauge theory. In this limit, the system has a global 1-form symmetry, $\cramped{A\to A+\lambda}$ with $\cramped{\mathrm{d}\lambda=0}$, called electric symmetry. This symmetry corresponds to electric strings forming closed loops, i.e. electric flux through closed surfaces is conserved. This theory has two phases, a confined phase at small $\beta$ (strong coupling), where the system is gapped and electric strings cost energy proportional to their length; and a deconfined phase at large $\beta$ (weak coupling), where the electric strings condense, the electric symmetry is spontaneously broken, and the system has a gapless photon excitation. Turning on a small $\kappa$ explicitly breaks the 1-form symmetry by introducing gapped electric point charges at which open electric strings end. Because the charges are strongly gapped, qualitatively speaking we expect the 1-form symmetry to re-emerge at low energies below the charge gap, thus allowing the deconfined phase to extend to finite $\kappa$.
@@
Strong coupling (confining) limit ($\beta=0$): In this limit the curvature of the gauge field $A$ is not penalized, and we can say that the magnetic monopoles are maximally proliferated. If we fix to unitary gauge ($\theta=\text{const}.$), the action becomes $\kappa\sum_{\ell}\cos(A_{\ell})$, reducing to a set of completely disconnected link variables. Thus the system is in a trivial phase for all $\kappa$, and there is no bulk phase transition on this line. Strong coupling expansion (in $\beta$) demonstrates that when $\kappa=0$ Wilson loops have an area law, indicating confinement of static charges by a tensionful electric string. This confinement extends to positive $\kappa$, though particle-anti-particle nucleation cuts off the area law of Wilson loops for large loops. Confinement may still be observed using the Fredenhagen-Marcu order parameter, however [27, 28].
Infinite Higgs coupling limit ($\kappa\to\infty$): In this limit we have the constraint $A=\mathrm{d}\theta$. This implies that the electric charge creation and annihilation operators, Eq.˜14, act as the identity, and the system can be described as an electric condensate. The bulk has no dynamics and is completely frozen, which can be seen by fixing to unitary gauge ($\theta=\text{const}.$), in which $A=0$ and the Higgs field is frozen. There is therefore no bulk phase transition along this line. For large but finite $\kappa$ the bulk action can be roughly understood as the proca-type action, Eq.˜23, describing a massive 1-form field, at least for large $\beta$.
This completes the general outline of the phase diagram in the vicinity of the edges in Fig.˜1. The only question that remains is how the two transitions at small $\kappa$ and large $\beta$ reach each other to separate the Coulomb phase (in which the electric and magnetic symmetries are emergent and spontaneously broken) from the Higgs-confined phase. Figure˜13 shows results from Monte Carlo simulations measuring the average Wilson plaquette and Wilson link, which fills in the remainder of the phase diagram schematically shown in Fig.˜1. The two transitions extend as first-order transition lines and meet at a triple point in the vicinity of $\beta\sim 1.0$ an $\kappa\sim 0.4$. A third first order line extends from the this triple point towards smaller $\beta$ and larger $\kappa$ which ends at a critical endpoint. We show evidence for this first-order line in Fig.˜13(c).